How many integers form 1 to 500 are divisible by 2, 3, 4, and5?


1
Expert's answer
2022-05-01T17:31:05-0400

500/2=250 numbers are divisible by 2, 

500/3= 166 numbers are divisible by 3. 

500/4= 125 numbers are divisible by 4, 

500/5= 100 numbers are divisible by 5. 


"3(4)(5)=60"

500/60=8 numbers are divisible by 2, 3, 4, and 5 simultaneously.